THE LAST LAP: HIGH-RESOLUTION INFRARED SPECTROSCOPY OF BUTADIENE-2.3-^{13}C_{2}.

\underline{NORMAN C. CRAIG} and MICHAEL C. MOORE, Department of Chemistry, Oberlin College, Oberlin, OH 44074; ROBERT L. SAMS, ESML Pacific Northwest National Laboratory, Richland, WA 99352; DONALD C. MCKEAN, School of Chemistry, University of Edinburgh, West Mains Road, Edinburgh EH9 3JJ, U. K..

This report on the high-resolution (0.002 cm^{−1}) infrared spectroscopy of butadiene-2,3-^{13}C_{2} is a culmination of the investigation of rotational constants of a series of isotopomers of butadiene (BDE). The goal is an equilibrium structure of butadiene with which to answer a question. Does the structure of butadiene reflect pi-electron delocalization? Ground state rotational constants are now available for BDE, BDE-2,3-d_{2},^{a} BDE-1,1-d_{2},^{b} BDE-trans,trans-1,4-d_{2}, BDE-cis,trans-1,4-d_{2}, BDE-cis,cis-1,4-d_{2},^{c} and BDE-1-^{13}C_{1}.^{d} BDE-2,3-^{13}C_{2} has been synthesized and its ground state rotational constants have been derived from the analysis of a C-type band arising from CH_{2} flapping at 907.1722 (2) cm^{−1}. These values, which are fit to 1184 ground state combination differences, include A = 1.3545028 (6), B = 0.1469410 (2), and C = 0.1325846 (2) cm^{−1}. Medium-resolution infrared and Raman spectroscopy have led to a complete assignment of vibrational fundamentals, which are useful in assessing a new force field for butadiene. The new force field depends on selectively scaled force constants in symmetry coordinate space. The final step in this project is obtaining good vibration-rotation alphas predicted by quantum chemical calculations for use in extracting equilibrium rotational constants and an equilibrium structure.

aN. C. Craig, J. L. Davis, K. A. Hanson, M. C. Moore, K. J. Weidenbaum, M. Lock, J. Mol. Struct. 2004, 695 − 696 , 59.

bW. Caminati, G. Grassi, A. Bauder, Chem. Phys. Lett. 1988, 148, 13.

cN. C. Craig, K. A. Hanson, R. W. Pierce, S. D. Saylor, R. L. Sams, J. Mol. Spectrosc. 2004, 228, 401.

dN. C. Craig, K. A. Hanson, M. C. Moore, R. L. Sams, J. Mol. Struct. 2005, in press.